OpenStudy (anonymous):

What is the similarity ratio of a prism with surface area 36 ft^2 to a similar prism with surface area 225 ft^2?

OpenStudy (anonymous):

@dlearner

OpenStudy (anonymous):

oops :P . no idea mate :)

OpenStudy (anonymous):

lol its all good, thanks anyway :)

OpenStudy (anonymous):

I think the formula is A1 / A2 = (l1 / l2)^2 where A1 and A2 are the areas and l1 / l2 is the similarity ratio

OpenStudy (anonymous):

how would i work that out exactly.. @TURITW

jimthompson5910 (jim_thompson5910):

hint: if the ratio of two sides of two figures is a/b, then the ratio of the two surface areas is (a/b)^2

OpenStudy (anonymous):

Since you know the areas, substitute into the formula and solve for l1 / l2

OpenStudy (anonymous):

Then, present your answer in the form of a ratio a : b.

OpenStudy (anonymous):

im not sure thats why i need your help...

jimthompson5910 (jim_thompson5910):

think of a/b as one variable, call it z if you want

OpenStudy (anonymous):

okay...

jimthompson5910 (jim_thompson5910):

your goal is to solve A1/A2 = z^2 for z

jimthompson5910 (jim_thompson5910):

A1 and A2 are the two surface areas

OpenStudy (anonymous):

woah i keep getting really big numbers

jimthompson5910 (jim_thompson5910):

don't square it, take the square root of both sides

OpenStudy (anonymous):

oooh ok

jimthompson5910 (jim_thompson5910):

ex: x^2 = 25 x = sqrt(5) or x = -sqrt(5) x = 5 or x = -5 notice how we used the square root to undo the squaring to isolate x

OpenStudy (anonymous):

2:5 ?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks for helping out!

jimthompson5910 (jim_thompson5910):

np